# Nonparametric estimation of the jump-size distribution for an M/G/1 storage system with Poisson sampling

### v3

@@ 3.3 Proofs of Propositions 3 and 4
We next construct bound the first series, i.e., the sum of $\mathrm{E}[Z_{1}(s)\overline{Z_{k}(s)}]$. An identical argument establishes the same bound for the last series with the term of $\mathrm{E}[Z_{k}(s)\overline{Z_{1}(s)}]$.
@@ 5 Concluding remarks
This work has presented a nonparametric method for the estimation of the job-size CDF in a storage system with periodic workload observations. Assuming that the distributions belong to a smoothness class satisfying $\lim_{s\to\infty}|\gamma(s)|s^{\eta}\leq C_{0}$, it was shown that the convergence rate of the risk is $O\left(n^{-\eta/(1+\eta)}\right)$. For some distributions in this class, the convergence rate is actually faster due to the faster decay of the bias term. For example, if $\lim_{s\to\infty}|\gamma(s)|\mathrm{e}^{-\omega s}\leq C$ for some $\omega>0$ (e.g., the Normal distribution truncated at zero), then an improved rate can be obtained. Specifically, applying the arguments used in the proof of Theorem 2 it can be shown that setting $h_{n}=L_{w}(\sqrt{n})\approx\log\sqrt{n}-\log\log\sqrt{n}$, where $L(\cdot)$ is the Lambert W function, yields a convergence rate of $O\left(L_{w}(\sqrt{n})^{2}/n\right)$. An open question that remains is whether the estimator is rate-optimal, in the sense that no other estimator based on the same workload observations achieves a faster rate of convergence. There is no reason to believe this is the case, but verifying that one can actually do better appears to be a very challenging problem due to the complex dependence structure of the data. As far as the author is aware, the estimator for $G$ presented here is the only nonparametric estimator with a rate guarantee for an M/G/1 system with workload observations. An additional issue is that the smoothness parameter $\eta$ of the underlying class of distributions is unknown, hence the truncation parameter $h_{n}$ must be selected heuristically in practice. This calls for development of an adaptive estimation technique that learns the smoothness parameter from the data (see [14] for such a scheme in the deconvolution setting).
@@
The approach presented here can also be applied to nonparametric estimation of the job-size density, assuming it exists and satisfies some regularity conditions. An interesting generalization is to extend the analysis to the setting of a Lévy-driven queue, where the input is some general subordinator and not necessarily a compound Poisson process (see [15]). The goal here would be to estimate the underlying Lévy measure of the input. The CF estimator in this case can be obtained directly from [31], which dealt with this general setting. However, inverting the net input CF to the Lévy measure is not straightforward. One possibility is to apply the inversion technique used in [22], which entails the estimation of the first and second derivatives of the CF. Another potential generalization is that of a Markov-modulated input process for which similar estimation equations can be constructed for the Poisson sampling scheme (see [4]). In the author’s view, these directions raise various interesting and challenging questions to be considered in future work.
@@ Acknowledgements
This work was supported by the Israel Science Foundation (ISF), grant no. 1361/23. The author is grateful to Alexander Goldenshluger and Michel Mandjes for many fruitful discussions and for their helpful feedback on drafts of this work.

### v4

@@ 3.3 Proofs of Propositions 3 and 4
We next construct a bound for the first series, i.e., the sum of $\mathrm{E}[Z_{1}(s)\overline{Z_{k}(s)}]$. An identical argument establishes the same bound for the last series with the term of $\mathrm{E}[Z_{k}(s)\overline{Z_{1}(s)}]$.
@@ 4 Unknown arrival rate
Figure 7 repeats the experiment for a system with lighter load. As opposed the case with known $\lambda$, here the estimation accuracy is actually lower for a system with a lower arrival rate. A possible explanation for this is that the estimation equation (39) becomes flatter, thus resulting in higher bias of the numerical integration step.
Figure 7: The estimators are based on a simulation of $n=40,000$ workload observations with sampling rate $\xi=0.5$. The arrival rate is $\lambda=0.2$ and the true CDF $G$ (black solid line) is a mixture of gamma distributions $\bm{\alpha}=(1.5,5)$, $\bm{\beta}=(0.8,10)$ and $\bm{p}=(0.4,0.6)$. The true CDF is compared with the estimator with the true $\lambda$ (with $h=8$) and the estimated $\widehat{\lambda}_{n}$ for $h\in\{4,6,8\}$.
@@ 5 Concluding remarks
This work has presented a nonparametric method for the estimation of the job-size CDF in a queue with periodic workload observations. Assuming that the distributions belong to a smoothness class satisfying $\lim_{s\to\infty}|\gamma(s)|s^{\eta}\leq C_{0}$, it was shown that the convergence rate of the risk is $O\left(n^{-\eta/(1+\eta)}\right)$. For some distributions in this class, the convergence rate is actually faster due to the faster decay of the bias term. For example, if $\lim_{s\to\infty}|\gamma(s)|\mathrm{e}^{-\omega s}\leq C$ for some $\omega>0$ (e.g., the Normal distribution truncated at zero), then an improved rate can be obtained. Specifically, applying the arguments used in the proof of Theorem 2 it can be shown that setting $h_{n}=L_{w}(\sqrt{n})\approx\log\sqrt{n}-\log\log\sqrt{n}$, where $L_{w}(\cdot)$ is the Lambert W function, yields a convergence rate of $O\left(L_{w}(\sqrt{n})^{2}/n\right)$. An open question that remains is whether the estimator is rate-optimal, in the sense that no other estimator based on the same workload observations achieves a faster rate of convergence. There is no reason to believe this is the case, but verifying that one can actually do better appears to be a very challenging problem due to the complex dependence structure of the data. As far as the author is aware, the estimator for $G$ presented here is the only nonparametric estimator with a rate guarantee for an M/G/1 system with workload observations. An additional issue is that the smoothness parameter $\eta$ of the underlying class of distributions is unknown, hence the truncation parameter $h_{n}$ must be selected heuristically in practice. This calls for development of an adaptive estimation technique that learns the smoothness parameter from the data (see [14] for such a scheme in the deconvolution setting).
@@
The approach presented here can also be applied to nonparametric estimation of the job-size density, assuming it exists and satisfies some regularity conditions. An interesting generalization is to extend the analysis to the setting of a Lévy-driven queue, where the input is some general subordinator and not necessarily a compound Poisson process (see [15]). The goal here would be to estimate the underlying Lévy measure of the input. The CF estimator in this case can be obtained directly from [31], which dealt with this general setting. However, inverting the net input CF to the Lévy measure is not straightforward. One possibility is to apply the inversion technique used in [22], which entails the estimation of the first and second derivatives of the CF. In the author’s view, these directions raise various interesting and challenging questions to be considered in future work.
Another extension to consider is that of a Markov-modulated input process for which similar estimation equations can be constructed for the Poisson sampling scheme. The main idea here is to utilize the framework of [4] that constructs an exponential (Wald type) martingale for the net-input process where the arrival rate is modulated by a CTMC with transition matrix $Q$ on a finite state space $\{\lambda_{1},\ldots,\lambda_{p}\}$. The CF of the net-input process can represented by the solution of $|Q+\mathrm{diag}({\bm{\varphi}})|=0$, where
$\varphi_{j}(s)=\lambda_{j}(\gamma(s)-1)-is\ ,j=1,\ldots,p\ .$
The methods presented in this work may be applied once the conditional workload distribution at Poisson sampling times is derived. Stochastic root finding will likely be necessary for the matrix inversion described above. An obvious obstacle is that knowledge of the arrival rates and transition matrix, or a reliable method for their estimation, is necessary.
@@ Acknowledgements
This work was supported by the Israel Science Foundation (ISF), grant no. 1361/23. The author thanks the anonymous reviewers for their valuable feedback. The author is grateful to Alexander Goldenshluger and Michel Mandjes for many fruitful discussions and for their helpful feedback on drafts of this work.